OpenStudy (anonymous):

Arrange these numbers from least to greatest: 3/5,..5/7,..2/9

OpenStudy (aum):

Make the denominator the same which would make it easier to compare the fractions. Find the least common denominator of 5, 7 and 9 and multiply top and bottom of each fraction by the least common denominator. Then compare.

OpenStudy (aum):

Or you can divide and express a couple of decimal places and then compare.

OpenStudy (anonymous):

Hint: \[\left\{\frac{3}{5},\frac{5}{7},\frac{2}{9}\right\}=\{0.6,0.714286,0.222222\} \]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Wait I don't get it

OpenStudy (aum):

\(\large \frac 35 = 0.6\) \(\large \frac 57 \approx 0.7\) \(\large \frac 29 \approx 0.2\) Arrange 0.6, 0.7 and 0.2 from least to greatest. Then you can arrange the corresponding fractions from least to greatest.